# Multi-Persona Thinking for Bias Mitigation in Large Language Models

### v1

@@ 4.3 Implementation Details
We evaluate our approach on four LLMs across different scales. For open-source models, we conduct comprehensive experiments on Llama-3.1-8B/70B-Instruct, evaluating both on BBQ and StereoSet. For closed-source models, we evaluate gpt-3.5-turbo with selected baselines on a randomly sampled BBQ subset (80 instances per category, 880 in total) due to resource constraints.
We set the maximum output length as 512 tokens for open-source models and 128 tokens for gpt-3.5-turbo. To ensure fairness of comparisons, we sample 3, 5, 10 and 15 independent outputs for SC. Following the implementation of previous work by  Du et al. (2023), we employ 3 agents in 3 rounds of debate for MAD. To match the computations, we also use 3 reasoning iterations for our MPT in the main experiments.
@@ 4.4 Main Results
Direct-Prompting Standard 0.6175 0.8976 0.7575 0.0839 0.0721 0.0780
Debias 0.6103 0.8801 0.7452 0.0676 0.0870 0.0773
Persona 0.5336 0.8657 0.6996 0.0880 0.0754 0.0817
Self-Consistency Standard 0.6448 0.9160 0.7804 0.0841 0.0677 0.0759
Debias 0.6461 0.9050 0.7755 0.0811 0.0793 0.0802
Persona 0.5574 0.9011 0.7292 0.1213 0.0704 0.0958
Re-Prompting Debias 0.8336 0.8077 0.8207 0.0387 0.0985 0.0686
Persona 0.7330 0.7055 0.7192 0.0359 0.1164 0.0761
MAD 0.8329 0.8311 0.8320 0.0357 0.1158 0.0758
MPT (ours) 0.9046 0.8755 0.8901 0.0275 0.0849 0.0562
@@
Direct-Prompting Standard 0.9273 0.8856 0.9064 0.0406 0.0115 0.0260
Debias 0.9282 0.7972 0.8627 0.0211 0.0109 0.0160
Persona 0.8396 0.7100 0.7748 0.0166 0.0108 0.0137
Self-Consistency Standard 0.9429 0.9219 0.9324 0.0424 0.0136 0.0280
Debias 0.9649 0.8623 0.9136 0.0223 0.0112 0.0168
Persona 0.9126 0.7543 0.8335 0.0156 0.0063 0.0109
Re-Prompting Debias 0.9833 0.8087 0.8960 0.0063 0.0341 0.0202
Persona 0.9686 0.7179 0.8432 0.0104 0.0373 0.0238
MAD 0.9886 0.8178 0.9032 0.0055 0.0127 0.0091
MPT (ours) 0.9846 0.8702 0.9274 0.0030 0.0077 0.0053
Table 3: Main results on the BBQ dataset. Accuracy ($\uparrow$) and Diff-bias score ($\downarrow$) are reported in ambiguous and disambiguated setups. Bold and underlined values denote the best and second-best micro-averaged results respectively.
Method Avg. Acc (95% CI) Avg. Diff-bias (95% CI)
MAD $8316\pm 0.0012 0.0$$0759\pm 0.0022$
MPT (ours) $\mathbf{0.8907\pm 0.0010}$ $\mathbf{0.0579\pm 0.0017}$
Paired t-test $t=169.68,p<0.001$ t=21.97,p<0.001t=-21.97,p<0.001
Table 4: Paired t-test of MPT vs. MAD on BBQ with Llama-3.1-8B-Instruct. The values are calculated from 5 independent runs. The results confirm the statistical significance of the improvements.
@@
   Standard 0.2360 0.1598
   Debias 0.2639 0.1407
   Persona 0.1927 0.1279
@@
   Standard 0.2289 0.1850
   Debias 0.2565 0.1681
   Persona 0.1545 0.1698
@@
   Debias 0.4561 0.1202
   Persona 0.3579 0.0993
MAD 0.2552 0.1089
MPT (ours) 0.5876 0.0635
@@
Table 3 presents the main results on the BBQ benchmark. With Llama-3.1-8B-Instruct, MPT achieves the highest average accuracy (89.01%) and the lowest average diff-bias score (0.0562), substantially outperforming all baselines. The improvement is particularly notable in the ambiguous cases (90.46% accuracy, 0.0275 diff-bias), where other methods are more prone to relying on stereotypes for making a prediction. On the contrary, MPT largely reduces the ambiguous diff-bias score, and leads to an improvement of 7 percentage points in accuracy. This indicates that MPT’s dialectical reasoning effectively prevents the model from defaulting to biased assumptions in uncertain contexts. On disambiguated examples, MPT maintains competitive accuracy (87.55%) while achieving substantially lower bias (0.0849) compared with re-prompting and MAD (significance tests are shown in Table 4). To ensure a fair comparison regarding computational cost, we provide an extended comparison with self-consistency (SC) in Appendix B (Table 12). The results show that even when sample runs of SC is increased to 15 (higher cost than MPT with $R=3$), MPT still achieves superior performance. This shows that the effectiveness of MPT derives from the qualitative reasoning mechanism rather than the mere quantitative computational scaling.
With the larger 70B model, we observe that model scaling improves accuracy and reduces bias in general, and that MPT continues to demonstrate superior performance. Although standard self-consistency achieves a slightly higher average accuracy (93.24%), our MPT’s accuracy is comparable (92.74%). However, MPT reduces the average diff-bias score by relative 81% from 0.0280 to 0.0053 compared to self-consistency (standard). The results show that even strong models exhibit residual stereotypes, MPT acts as a last-mile safety mechanism that removes the remaining biases where scaling alone fails to resolve.
@@
The results on StereoSet are shown in Table 5. Since StereoSet contexts are always ambiguous, performance on this benchmark is a strong indicator of a model’s reliance on stereotypes. As we see, MPT again achieves the best results in both accuracy (58.76%) and diff-bias score (0.0635), significantly outperforming all baselines. This represents a 28% relative improvement in accuracy and a 36% relative reduction in bias compared with the second-best methods (re-prompting with explicit-debiasing for accuracy and persona-based re-prompting for the diff-bias score). This confirms that multi-persona reasoning is particularly effective in uninformative contexts where models must resist defaulting to stereotypical associations. Detailed breakdowns by test type (Table 13 in Appendix B) show consistent gains across both word-level and sentence-level subsets.
@@ 4.5 Analysis
Method amb disamb average
Accuracy $\uparrow$
@@
   Debias 0.9636 0.5525 0.7580
MAD 0.9292 0.4116 0.6704
MPT (ours) 0.8506 0.6492 0.7499
Diff-bias Score $\downarrow$
Direct-Prompting
   Standard 0.0636 0.0517 0.0577
@@
   Debias 0.0273 0.0132 0.0203
MAD 0.0068 0.0528 0.0298
MPT (ours) 0.0115 0.0166 0.0141
Table 6: Results of gpt-3.5-turbo on BBQ (80 instances per category, 880 in total).
Method Avg. Acc. $\uparrow$ Avg. Diff-bias $\downarrow$
@@
   w/o Neutral Persona 0.7775 0.0748
   w/ Neutral Persona 0.8901 0.0562
@@
Figure 2: Performance of Llama-3.1-8B-Instruct on BBQ with different numbers of MPT’s dialectical reasoning iterations ($R$). Top: accuracy $\uparrow$. Bottom: diff-bias score $\downarrow$.
Figure 3: Effect of applying self-consistency to MPT on BBQ using Llama-3.1-8B-Instruct. Both accuracy $\uparrow$ (left) and diff-bias $\downarrow$ (right) are improved notably.
@@
In this analysis, we test gpt-3.5-turbo on BBQ with a selection of representative baselines. The results in Table 6 highlight MPT’s capability to effectively balance accuracy and fairness. It achieves the lowest overall diff-bias score (0.0141), a 30% relative improvement over the second-best method. Although re-prompting scores are marginally higher on accuracy (75.8% vs. 74.99%), it comes at the cost of higher bias score (0.0203 vs. 0.0141). This trade-off highlights MPT’s strength in balancing accuracy and fairness in different model families.
@@
We analyze the critical role of the neutral general persona during MPT’s iterative reasoning phase. As shown in Table 7, the ablation study demonstrates a clear necessity of this component. Excluding the neutral persona results in significant performance degradation, where average accuracy drops to 77.75% and diff-bias score increases to 0.0748. These results align with the findings from prior work Borah and Mihalcea (2024), where prompting with binary contrasting personas amplifies bias in multi-agent setups and leads to polarization. This shows that the neutral persona helps the model balance the different perspectives effectively to achieve higher accuracy and lower bias score.
@@
Figure 2 illustrates the impact of reasoning rounds ($R$) on performance. At $R=0$, where conclusions are drawn directly from the initial generations without any dialectical reasoning, the model relies heavily on internal stereotypical assumptions, resulting in lower accuracy and higher bias score in ambiguous contexts. The figure shows a substantial leap in accuracy and a reduction in diff-bias score in ambiguous setups at $R=1$, and the performance converges rapidly beyond $R=2$. This demonstrates that a few rounds of multi-persona interactions are sufficient for the model to converge to debiased answers.
Regarding disambiguated contexts, MPT maintains high accuracy and even improves slightly when $R>0$, confirming that the reasoning capability is preserved. While there is a slight increase in the diff-bias score for disambiguated cases, suggesting a side effect of the model’s sensitivity to diverse perspectives. This is outweighed by the gains in ambiguous contexts.
@@
To test whether MPT can be combined with other prompting methods, we integrated it with self-consistency by sampling 5 independent MPT outcomes and majority voting the final answer. As shown in Figure 3, this combination consistently improves the performance on BBQ with Llama-3.1-8B-Instruct. The average accuracy increases from 89.01% to 92.32%, while the average diff-bias score decreases from 0.0562 to 0.0546. This indicates that MPT is not a monolithic solution but a robust and composable framework that can be effectively combined with other techniques to further enhance reliability and fairness.
@@ 5 Conclusion
This paper introduces Multi-Persona Thinking (MPT), a novel inference-time framework for mitigating social bias in LLMs. By engaging models in an iterative, dialectical process across different social identities, MPT transforms the potential weakness of persona-assignment into a strength for bias mitigation, and establishes a robust process for self-debiasing that simultaneously reduces bias and preserves, and often enhances, logical reasoning.
Our experiments and significance tests confirm that MPT consistently outperforms strong baselines by successfully overcoming the common fairness-performance trade-off. We further demonstrate that the framework is compatible with other techniques for additional performance improvement. Ultimately, MPT can be a practical tool that helps make LLMs fairer and more reliable.
@@ 6 Limitations
While MPT successfully demonstrates a method for balancing fairness and reasoning, it has several limitations.
First, although MPT is efficient compared to scaling sample sizes (e.g., SC with $k=15$), it still needs a higher inference cost and latency than standard direct prompting due to its multi-turn, multi-persona nature. However, this trade-off is adjustable: users can flexibly control the number of reasoning rounds based on latency requirements, enabling practical deployment in both real-time and offline scenarios.
Second, our current implementation uses predefined personas derived from dataset metadata (e.g., male vs. female). We acknowledge that this approach assumes social identities are stable and easily delineated, which is a simplification. Real-world identities are often complex, non-binary, and intersectional. While metadata-based personas serve as a necessary starting point for controlled evaluation on existing benchmarks, they may not fully capture the complexity of social dynamics. Moreover, our framework is flexible. For real-world queries lacking metadata, MPT does not strictly require specific social identity personas, as personas can be abstract or combined to more comprehensively represent diverse perspectives (e.g., an observer sensitive to gender). Future work could explore dynamic persona generation to address these intersectional and implicit identity attributes.
Third, our current evaluation focuses on multiple-choice tasks. Although open-ended generation is more complex, the core idea of using diverse personas to expose and resolve bias is theoretically transferable. However, empirical validation in such settings faces two significant hurdles. The first one is the shortage of unbiased evaluation metrics: Open-ended generation lacks a universally accepted, unbiased gold standard for evaluation. Current automated metrics (e.g., LLM-as-a-judge) may contain their own inherent biases, which could lead to biased judgments. The second one is the scarcity of controlled benchmarks: The lack of datasets that match the rigor of BBQ, which rigorously controls variables to simultaneously evaluate reasoning capabilities and social biases. Without such controlled settings, it becomes challenging to measure whether an improvement in bias comes at the cost of reasoning in open-ended generation.
@@ Appendix A Dataset Details
Category Number of examples
@@
Table 8: Total number of examples within each of BBQ’s categories.
@@
Word-level
  gender 771
  race 2976
  religion 247
  profession 2398
Sentence-level
  gender 751
  race 2947
  religion 241
  profession 2435
Total 12,766
@@
Tables 8 and 9 provide detailed statistics for the BBQ and StereoSet datasets. BBQ covers 11 social bias categories with over 58k instances, where around 46% are intersectional biases. StereoSet contains around 12.8k examples of word-level and sentence-level tasks. Examples of each task in gender bias are illustrated in Tables 10 and 11.
Appendix B More Experimental Results

### v3

@@ 4.3 Implementation Details
We evaluate our approach on four LLMs of different scales. For open-source models, we conduct comprehensive experiments on Llama-3.1-8B/70B-Instruct and Qwen-2.5-7B-Instruct, evaluating them both on BBQ and StereoSet. For closed-source models, we evaluate GPT-3.5-Turbo with selected baselines on a randomly sampled BBQ subset (80 instances per category, 880 in total) due to resource constraints.
We set the maximum output length to 512 tokens for open-source models and 128 tokens for the closed-source model. To ensure a fair comparison, we evaluate self-consistency with 3, 5, 10, and 15 samples. Following  Du et al. (2023), we employ three agents in three rounds of debate for MAD. To match the computational cost, we also use three reasoning iterations for our MPT in the main experiments.
@@ 4.4 Main Results
Direct-Prompting Standard 0.6005 0.8995 0.7499 0.0782 0.0706 0.0744
Debias 0.6222 0.8806 0.7514 0.0703 0.0850 0.0777
Persona 0.5362 0.8604 0.6983 0.0933 0.0734 0.0833
Self-Consistency Standard 0.6288 0.9195 0.7741 0.0882 0.0637 0.0759
Debias 0.6592 0.9079 0.7835 0.0823 0.0766 0.0794
Persona 0.5607 0.9037 0.7322 0.1288 0.0646 0.0967
Re-Prompting Debias 0.8324 0.8097 0.8210 0.0369 0.1022 0.0696
Persona 0.7313 0.7082 0.7198 0.0389 0.1138 0.0763
MAD 0.8332 0.8300 0.8316 0.0337 0.1181 0.0759
MPT (ours) 0.9054 0.8761 0.8907 0.0279 0.0880 0.0579
@@
Direct-Prompting Standard 0.9188 0.8675 0.8931 0.0407 0.0208 0.0307
Debias 0.8976 0.7666 0.8321 0.0179 0.0100 0.0140
Persona 0.7982 0.6829 0.7405 0.0148 0.0124 0.0136
Self-Consistency Standard 0.9401 0.9138 0.9269 0.0430 0.0173 0.0301
Debias 0.9556 0.8348 0.8952 0.0222 0.0096 0.0159
Persona 0.8790 0.7152 0.7971 0.0152 0.0055 0.0103
Re-Prompting Debias 0.9853 0.7932 0.8892 0.0055 0.0319 0.0187
Persona 0.9662 0.7189 0.8426 0.0079 0.0332 0.0205
MAD 0.9884 0.8190 0.9037 0.0050 0.0135 0.0092
MPT (ours) 0.9855 0.8711 0.9283 0.0034 0.0072 0.0053
Table 3: Main results on the BBQ dataset. Accuracy ($\uparrow$) and Diff-bias score ($\downarrow$) are reported in ambiguous and disambiguated setups. Values are averaged over five independent runs. Bold and underlined values denote the best and second-best results respectively.
@@
   Standard 0.2383 0.1654
   Debias 0.2473 0.1317
   Persona 0.1954 0.1239
@@
   Standard 0.2293 0.1900
   Debias 0.2338 0.1633
   Persona 0.1585 0.1636
@@
   Debias 0.4664 0.0974
   Persona 0.3532 0.0888
MAD 0.2706 0.1188
MPT (ours) 0.6073 0.0505
Qwen-2.5-7B-Instruct
Direct-Prompting
   Standard 0.5025 0.1981
   Debias 0.5849 0.1665
   Persona 0.5283 0.1873
Self-Consistency
   Standard 0.5020 0.1998
   Debias 0.5868 0.1673
   Persona 0.5284 0.1881
Re-Prompting
   Debias 0.7019 0.1007
   Persona 0.6668 0.1110
MAD 0.5268 0.1795
MPT (ours) 0.7312 0.0921
@@
Table 3 presents the main results on the BBQ benchmark. With Llama-3.1-8B-Instruct, MPT achieves the highest average accuracy (89.07%) and the lowest average diff-bias score (0.0579). The significance tests in Appendix B.1 confirm that MPT statistically outperforms all baselines with the 8B model. The improvement is particularly notable in the ambiguous cases (90.54% accuracy, 0.0279 diff-bias), where other methods are more prone to relying on stereotypes. In contrast, MPT largely reduces the ambiguous diff-bias score and leads to an improvement of 7 percentage points in accuracy. This indicates that MPT’s dialectical reasoning effectively prevents the model from defaulting to biased assumptions in uncertain contexts. On disambiguated examples, MPT maintains competitive accuracy (87.61%) while achieving substantially lower bias (0.0880) compared to re-prompting and MAD. To ensure a fair comparison in terms of computational cost, we provide an extended comparison with self-consistency in the Appendix C. The results show that even when SC is scaled up to 15 samples (at a higher cost than MPT with $R=3$), MPT still achieves superior performance. This indicates that the effectiveness of MPT derives from the qualitative reasoning mechanism rather than the mere increases in computational scale.
With the larger 70B model, we observe that model scaling improves accuracy and reduces bias score in general. Based on the average results of five independent runs, MPT achieves the highest overall performance in both metrics, confirming its overall superiority on the 70B scale. Specifically, MPT achieves an average accuracy of 92.83%, which is slightly higher than that of the strongest baseline (standard SC with 92.69%). Although significance tests (Appendix B.2) show that this improvement is not statistically significant, MPT maintains a comparable top-tier performance level in accuracy. However, MPT reduces the average diff-bias score by a relative 82% from 0.0301 to 0.0053.
@@
The results on StereoSet are shown in Table 4. Since all contexts in StereoSet are “ambiguous”, performance on this benchmark is a strong indicator of a model’s reliance on stereotypes. On Llama-3.1-8B-Instruct, MPT achieves the best results in both accuracy (60.73%) and diff-bias score (0.0505), significantly outperforming all baselines. This represents a 30% relative improvement in accuracy and a 43% relative reduction in bias compared with the second-best methods (re-prompting with explicit-debiasing for accuracy and persona-based re-prompting for the diff-bias score). The results on Qwen-2.5-7B-Instruct show the same trend, with MPT consistently outperforming the baselines (statistical tests in Appendix B.3). This indicates that multi-persona reasoning is particularly effective in uninformative contexts where models rely on stereotypical assumptions. Detailed breakdowns by test type are shown in Appendix D.
@@ 4.5 Analysis
[…]
GPT-3.5-Turbo
@@
   Debias 0.7580 0.0203
MPT (ours) 0.7499 0.0141
Qwen-2.5-7B-Instruct
Self-Consistency
   Debias 0.8673 0.0686
@@
   Persona 0.7440 0.0262
MPT (ours) 0.8913 0.0266
Table 5: Results of best methods with GPT-3.5-Turbo and Qwen-2.5-7B-Instruct on BBQ subsets.
[…]
@@
   w/o Neutral 0.7775 0.0748
   w/ Neutral 0.8901 0.0562
@@
Figure 2: Performance of Llama-3.1-8B-Instruct on BBQ with different numbers of MPT’s dialectical reasoning iterations ($R$).
Figure 3: Effect of applying self-consistency to MPT on BBQ using Llama-3.1-8B-Instruct. Both accuracy (left) and diff-bias (right) are improved notably.
@@
We further evaluate MPT’s performance on GPT-3.5-Turbo and Qwen-2.5-7B-Instruct with subsets of the BBQ dataset. The results of the best-performing methods for each model are summarized in Table 5 (comprehensive results provided in Appendix D Table 20). On GPT-3.5-Turbo, MPT achieves the lowest overall diff-bias score (0.0141), a relative improvement of 30% over the second-best method. Although re-prompting scores are marginally higher on accuracy (75.8% vs. 74.99%), it comes at the cost of a higher bias score (0.0203 vs. 0.0141). For Qwen-2.5-7B-Instruct, although re-prompting achieves a slightly lower diff-bias score, paired t-tests (Appendix B.4) indicate that this difference is not statistically significant. In contrast, its accuracy is much lower than that of MPT. The trade-offs highlight the strength of MPT in balancing accuracy and bias score across different model families.
@@
We analyze the role of the neutral general persona in MPT’s iterative reasoning process. Table 6 shows that this component is important for both accuracy and diff-bias score. Excluding the neutral persona results in significant performance degradation, where average accuracy drops to 77.75% and diff-bias score increases to 0.0748. These results align with findings from previous work (Borah and Mihalcea, 2024), which report that using only binary contrasting personas in multi-agent settings can amplify bias and lead to polarization. Our results suggest that the neutral persona helps maintain balance between perspectives and supports better overall performance.
@@
Figure 2 illustrates the impact of reasoning rounds ($R$) on performance. At $R=0$, where conclusions are drawn directly from the initial generations without any dialectical reasoning, the model relies heavily on internal stereotypical assumptions, resulting in lower accuracy and a higher bias score in ambiguous contexts. The figure shows a substantial leap in accuracy and a reduction in diff-bias score in ambiguous setups at $R=1$, and the performance converges rapidly beyond $R=2$. This demonstrates that a few rounds of multi-persona interaction are sufficient for the model to converge to debiased answers.
Regarding disambiguated contexts, MPT maintains high accuracy and even improves slightly when $R>0$, suggesting that reasoning ability is not compromised. The diff-bias score shows a slight increase in disambiguated cases, which may reflect a side effect of the model’s sensitivity to diverse perspectives. However, this increase is outweighed by the gains in ambiguous contexts.
@@
To test whether MPT can be combined with other prompting methods, we integrate it with self-consistency by sampling five independent MPT outputs and using majority voting for the final answer. As shown in Figure 3, this combination consistently improves performance on BBQ with Llama-3.1-8B-Instruct. The average accuracy increases from 89.01% to 92.32%, while the average diff-bias score decreases from 0.0562 to 0.0546. The improvement is observed in both ambiguous and disambiguated subsets. The results indicate that MPT is a robust and flexible framework that can be effectively combined with other techniques to further improve both reliability and fairness.
@@ 5 Conclusion
In this paper, we introduce Multi-Persona Thinking, an inference-time framework for mitigating social bias in LLMs. MPT engages the model in an iterative reasoning process across different social identities, turning persona assignment from a potential weakness into a useful mechanism for bias mitigation. Establishes a robust self-debiasing process that simultaneously reduces bias while preserving, and often enhancing, logical reasoning.
Our experiments show that MPT consistently outperforms strong baselines and alleviates the common trade-off between fairness and performance. Our analysis demonstrates that MPT can be effectively combined with other techniques for further improvement. In general, MPT is a practical and effective approach to making LLMs more fair and reliable.
@@ 6 Limitations
Although MPT provides an effective way to balance fairness and reasoning, it has several limitations.
First, although MPT is more efficient than simply increasing the number of samples (e.g., SC with 15 sampled runs), it still incurs higher inference cost and latency than standard direct prompting because it relies on multiple turns and multiple personas. However, this trade-off is adjustable. Users can control the number of reasoning rounds based on their latency constraints, which makes MPT applicable to both real-time and offline settings.
Second, our current implementation uses predefined personas derived from dataset metadata (e.g., male vs. female). This design assumes that social identities are stable and clearly separable, which is a simplification. In reality, identities are often complex, non-binary, and intersectional. Although metadata-based personas provide a practical starting point for controlled evaluation on existing benchmarks, they cannot fully capture the complexity of real social contexts. Moreover, our framework is flexible. For real-world queries without explicit metadata, MPT does not require fixed social-identity personas. Instead, personas can be abstract, combined, or designed to reflect broader perspectives (e.g., an observer sensitive to gender bias). Future work could explore the generation of dynamic persona to better handle implicit and intersectional identities.
Third, our current evaluation focuses on multiple-choice tasks. Although open-ended generation is more complex, the core idea of using diverse personas to expose and resolve bias is transferable, in principle. However, empirical validation in such settings faces two major challenges. The first is the lack of unbiased evaluation metrics. Open-ended generation lacks a universally accepted, unbiased gold standard for evaluation. Current automated metrics such as LLM-as-a-judge may themselves introduce bias, which could lead to biased judgments. The second is the scarcity of controlled benchmarks. Unlike BBQ, which carefully controls variables to evaluate both reasoning ability and social bias, existing open-ended benchmarks rarely provide the same level of assessment. As a result, it is difficult to determine whether a reduction in bias comes at the cost of reasoning quality in open-ended settings.
7 Acknowledgments
We thank the reviewers and chairs for their efforts. The research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), the Amii Fellow Program, the Canada CIFAR AI Chair Program, a donation from DeepMind, and the Digital Research Alliance of Canada (alliancecan.ca).
@@ Appendix A Dataset Details
Table 7 and Table 8 provide detailed statistics for the BBQ and StereoSet datasets. BBQ covers 11 categories of social bias with more than 58k instances, where around 46% are intersectional biases. StereoSet contains around 12.8k instances of both word- and sentence-level tasks. Table 9 shows examples of contrasting pair of persona from each category. Examples of each task in gender bias are illustrated in Tables 10 and 11.
Category Number of instances
@@
Table 7: Total number of examples within each of BBQ’s categories. Each category contains the same number of ambiguous and disambiguated tasks.
@@
Word-level Sentence-level Total
Gender 771 751 1,522
Race 2,976 2,947 5,923
Religion 247 241 488
Profession 2,398 2,435 4,833
Total 6,392 6,374 12,766
@@
Category Contrasting persona example
age grandfather vs. grandson
disability status person in a wheelchair vs. person who bikes in to work
gender identity man vs. woman
nationality Indian vs. Japanese
physical appearance fat vs. skinny
race/ethnicity Black vs. White
religion Muslim vs. Hindu
sexual orientation gay vs. straight
socio-economic status wealthy vs. poor
race by gender Black man vs. White woman
race by SES Black wealthy vs. Latino poor
Table 9: Contrasting persona pairs instantiated in MPT for each BBQ category.
@@
Appendix B Statistical Tests
To further validate the effectiveness of MPT, we conduct significance tests against the strongest baselines. All values are collected from five independent runs, and statistical significance is assessed using paired t-tests..
B.1 Llama-3.1-8B-Instruct on BBQ
Method Avg. Acc (95% CI) Avg. Diff-bias (95% CI)
MAD $8316\pm 0.0012 0.0$$0759\pm 0.0022$
MPT (ours) $\mathbf{0.8907\pm 0.0010}$ $\mathbf{0.0579\pm 0.0017}$
Paired t-test $t=169.68,p<0.001$ t=21.97,p<0.001t=-21.97,p<0.001
Table 12: Paired t-test of MPT vs. MAD on BBQ with Llama-3.1-8B-Instruct.
Method Avg. Acc (95% CI) Avg. Diff-bias (95% CI)
Re-prompting (Debias) $8210\pm 0.0016 0.0$$0696\pm 0.0014$
MPT (ours) $\mathbf{0.8907\pm 0.0010}$ $\mathbf{0.0579\pm 0.0017}$
Paired t-test $t=160.96,p<0.001$ t=18.19,p<0.001t=-18.19,p<0.001
Table 13: Paired t-test of MPT vs. Re-prompting (Debias) on BBQ with Llama-3.1-8B-Instruct.
Table 12 and Table 13 show that on BBQ with Llama-3.1-8B-Instruct, MPT outperforms the second-best baseline on both accuracy (MAD) and diff-bias score (re-prompting).
B.2 Llama-3.1-70B-Instruct on BBQ
Method Avg. Acc (95% CI) Avg. Diff-bias (95% CI)
SC (Standard) $9269\pm 0.0040 0.0$$0301\pm 0.0023$
MPT (ours) $\mathbf{0.9283\pm 0.0007}$ $\mathbf{0.0053\pm 0.0009}$
Paired t-test $t=0.83,p=0.451$ t=27.97,p<0.001t=-27.97,p<0.001
Table 14: Paired t-test of MPT vs. SC (Standard) on BBQ with Llama-3.1-70B-Instruct.
Method Avg. Acc (95% CI) Avg. Diff-bias (95% CI)
MAD $9037\pm 0.0004 0.0$$0092\pm 0.0001$
MPT (ours) $\mathbf{0.9283\pm 0.0007}$ $\mathbf{0.0053\pm 0.0009}$
Paired t-test $t=171.81,p<0.001$ t=12.42,p<0.001t=-12.42,p<0.001
Table 15: Paired t-test of MPT vs. MAD on BBQ with Llama-3.1-70B-Instruct.
For the Llama-3.1-70B-Instruct model, Table 14 indicates that while MPT does not show a statistically significant improvement in accuracy over self-consistency ($p=0.451$), it remains a top-tier performance in terms of accuracy. Crucially, the results in Table 15 demonstrate the superiority of MPT on diff-bias score.
B.3 Qwen-2.5-7B-Instruct on StereoSet
Method Avg. Acc (95% CI) Avg. Diff-bias (95% CI)
Re-prompting (Debias) $0.7019\pm 0.0019$ $0.1007\pm 0.0018$
MPT (ours) $\mathbf{0.7312\pm 0.0015}$ $\mathbf{0.0921\pm 0.0029}$
Paired t-test t=26.03,p<0.001t=26.03,\ p<0.001 t=7.65,p=0.0016t=-7.65,\ p=0.0016
Table 16: Paired t-test of MPT vs. Re-prompting (Debias) on StereoSet with Qwen-2.5-7B-Instruct.
On the StereoSet dataset, Table 16 shows the statistical superiority of MPT over re-prompting (Debias) for Qwen-2.5-7B-Instruct.
B.4 Qwen-2.5-7B-Instruct on BBQ Subset
Method Avg. Acc (95% CI) Avg. Diff-bias (95% CI)
Re-prompting (Persona) $7440\pm 0.0044 0.0$$0262\pm 0.0052$
MPT (ours) $\mathbf{0.8913\pm 0.0025}$ $\mathbf{0.0266\pm 0.0159}$
Paired t-test t=61.25,p<0.001t=61.25,\ p<0.001 t=0.08,p=0.9422t=0.08,\ p=0.9422
Table 17: Paired t-test of MPT vs. Re-prompting (Persona) on BBQ subset with Qwen-2.5-7B-Instruct.
Regarding the BBQ subset, results in Table 17 show that while re-prompting (Persona) achieves a slightly lower diff-bias score, the difference is not statistically significant compared to MPT ($p=0.9422$). Crucially, MPT achieves this comparable level of diff-bias score with a significant improvement in accuracy.
Appendix C Cost Match Comparison